OpenStudy (anonymous):

Compute the permutation. 7 P 2 2 42 49

OpenStudy (anonymous):

@Marki

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aaronandyson):

http://www.regentsprep.org/regents/math/algebra/APR2/Lperm.htm this link might be helpful

OpenStudy (anonymous):

49?

OpenStudy (aaronandyson):

so that is \[7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1 \over 5 \times 4 \times 3 \times 2 \times 1\]

OpenStudy (aaronandyson):

as we took \[7! \over (7- 2)!\]

OpenStudy (anonymous):

it turnes out around 600?

OpenStudy (aaronandyson):

WHAT1?

OpenStudy (aaronandyson):

CANCEL THE COMMON NUMBER UP THERE

OpenStudy (anonymous):

wait actaully 42

OpenStudy (anonymous):

i accidentally did times instead of divide lol

OpenStudy (anonymous):

thnx:))

OpenStudy (aaronandyson):

np :)